OpenStudy (anonymous):

Write an equation for the line that contains the point (-2, 9) and is parallel to the line y = 2x + 6. A. y = -(1/2)x + 8 B. y = 2x – 7 C. y = 2x + 11 D. y = 2x + 13

OpenStudy (anonymous):

If the line is parallel, it has the same slope. So we can just plug in the given x and y values into the original equation. 9=2(-2)+b 9=-4+b 13=b So the equation would be y=2x+13 so the answer is D.

OpenStudy (anonymous):

Thank you for explaining!

OpenStudy (anonymous):

No problem :)